# GenoCraft: A Comprehensive, User-Friendly Web-Based Platform for High-Throughput Omics Data Analysis and Visualization

### v2

@@
Yingzhou Lu1,2, Minjie Shen2, Ling Yue3, Chenhao Li4, Fan Meng2, Xiao Wang5,
@@ 3 Results
3.1 Case Study: Bulk RNA-seq Analysis of GSE99611
GSE99611 (https://ftp.ncbi.nlm.nih.gov/geo/series/GSE99nnn/GSE99611/matrix/) is a bulk RNA-sequencing dataset available on the Gene Expression Omnibus (GEO) database. It contains gene expression information generated from a specific biological study [BTW+10]. Our aim in this case study is to elucidate the underlying biological insights from this dataset by using a comprehensive RNA-seq analysis pipeline.
After obtaining raw read counts, we proceed to the normalization of these counts and differential expression analysis. These tasks are performed using the t-test. These tools not only allow for normalization - accounting for variations in sequencing depth and gene length across samples - but also provide a statistical framework to identify differentially expressed genes. Ultimately, this analysis yields significant genes whose expression levels vary noticeably across different conditions, providing critical insights into the biological phenomena under investigation in the GSE99611 dataset.

### v3

@@
Yingzhou Lu1,2, Minjie Shen2, Ling Yue3, Chenhao Li4, Lulu Chen2, Xiao Wang5,
@@ 3 Results
3.1 Case Study: GSE152418
In this case study, we applied GenoCraft’s RNA-seq analysis pipeline to the GSE152418 dataset from the Gene Expression Omnibus (GEO). This dataset explores the immune response in patients with COVID-19 by analyzing gene expression in peripheral blood mononuclear cells (PBMCs) from both COVID-19 patients and healthy convalescent controls.
The raw counts for gene expression were retrieved from GSE152418, which consists of samples from COVID-19 patients and convalescent controls. The dataset was structured with gene identifiers (ENSEMBL IDs) as rows and samples as columns, including 17 COVID-19 patients and their respective control counterparts. Initial quality control was performed by removing samples and genes with low counts, ensuring the data was suitable for downstream analysis.
To account for differences in sequencing depth and library sizes across samples, we applied Counts Per Million (CPM) normalization to the raw read counts. This normalization technique adjusts the gene expression values, making them comparable across the dataset. CPM is particularly useful when working with RNA-seq data, as it mitigates the effects of varying sequencing depths across samples.
We proceeded with the differential expression analysis, comparing the COVID-19 group against the convalescent control group. Using GenoCraft’s integrated t-test, we calculated log fold changes (LFC) between the two groups to identify genes that were significantly upregulated or downregulated in COVID-19 patients.
Significant genes were identified based on their fold change and statistical significance. These genes represent key players in the immune response to COVID-19 and provide valuable insights into the molecular mechanisms driving disease pathology.
The differentially expressed genes (DEGs) were further subjected to pathway enrichment analysis using the well-known databases. By mapping DEGs to known immune response pathways, we identified critical pathways activated in COVID-19 patients. These findings revealed potential biomarkers and therapeutic targets relevant to the immune dysregulation observed in COVID-19.
GenoCraft’s suite of visualization tools provided intuitive representations of the analysis results. Heatmaps illustrated the expression patterns of the most significant genes across samples, while volcano plots highlighted genes with the highest fold changes. These visualizations offered a clear and comprehensive overview of the data, facilitating interpretation and aiding in the identification of biologically relevant gene expression changes.
The analysis of the GSE152418 dataset using GenoCraft underscores the platform’s capabilities in handling large-scale RNA-seq datasets. By identifying differentially expressed genes and key immune pathways, GenoCraft provided actionable insights into the biological response to COVID-19, potentially informing future research and therapeutic development.